OpenStudy (anonymous):

7. Assuming g(x) = x2 + 3, what is the change made to g(x) if the parent function is f(x) = x2? A. Vertical translation of 3 units up B. Vertical translation of 3 units down C. Horizontal translation of 3 units to the right D. Horizontal translation of 3 units to the left

OpenStudy (mathmale):

One way to approach this problem would be for you to graph both\[g(x) = x^2 + 3~and~f(x)=x^2.\]

OpenStudy (mathmale):

Starting with the graph of f(x), how do you translate (move) that graph so that it becomes the graph of g(x)?

OpenStudy (anonymous):

i graphed itand i got D @mathmale